Spot the Similarity and Difference!


Below are three expressions. Observe them and spot their similarities and


differences. (a) 3-281


(b) y = 4


(c)2 à ¢ ‰ ¥ 32


as


Questions:


1. What is the similarity of (a), (b), and (c)?


2. What is the similarity of (a) and (c)?


3. How do (a) and (c) differ from (b)?


4. How are (a) and (b) similar to each other?


5. How do (a) and (b) differ from (c)? 6. Can you recall which among them is an exponential function?


7. Which among the three is/are exponential expression/s?


1
Expert's answer
2021-10-18T13:05:46-0400

a) 3x2=813^{x-2}=81 ; b) y=4xy=4^x ; c) 2x322^x\ge 32


1.

all are expressions with a variable used as an exponent


2.

(a) and (c) both involve one variable only


3.

(a) and (c) both involve one variable only while (b) involves the relationship between two variables


4.

(a) and (b) both use an equal sign


5.

(a) and (b) both use an equal sign while (c) contains an inequality symbol


6.

b is an exponential function since it shows the relation between dependent variable y and independent variable x


7.

All three are exponential expressions